OpenStudy (anonymous):

The ages of a group of women are approximately normally distributed with a mean of 48 years and a standard deviation of 6 years. One woman is randomly selected from the group, and her age is observed. A. Find the probability that her age will fall between 54 and 60 years is_____. B. Find the probability that her age will fall between 47 and 52 years is_____. C. Find the probability that her age will be less than 35 is_____. D. Find the probability that her age will exceed 40 years is_____.

OpenStudy (amistre64):

are you spose to use an exact result, or an empirical rule approximation?

OpenStudy (amistre64):

also, are you using a ti83, or stat program, or do you have to use the tables in the book?

OpenStudy (anonymous):

Tables in the book

OpenStudy (amistre64):

tables bite .... different authors have different formats that they feel are better :/ in any case, the process is simple enough if you can read the table

OpenStudy (amistre64):

find a z score, look it up in the table, work the results\[Z=\frac{X-mean}{sd}\]

OpenStudy (amistre64):

since your mean = 48, and sd = 6 finding the z scores for the given info is simple enough

OpenStudy (amistre64):

A. Find the probability that her age will fall between 54 and 60 years z1 = (54-48)/6 ; z2 = (60-48)/6 do you know how to read your tables?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

then the rest of it is working out the values from the table to address the given situations.

OpenStudy (anonymous):

ok thanks for you help

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

By the way the last two quetions, would you use 35 or 34 if it is less than, and 36 if it is greater tahn?

OpenStudy (amistre64):

C. Find the probability that her age will be less than 35 is_____. x = 35, then find the area to the left of z D. Find the probability that her age will exceed 40 years is_____. x = 40, then find the area to the right of z